Roosevelt, Inc. produces the following income statement and balance sheet for the year. Net sales of the firm are 3800million. Depreciation is 360million. Earnings before income and taxes is 1780million and net income is 832million. Interest expense is 500million. Current liabilities, long term debt and shareholder’s equity are 800, 2000 and 4000 million, respectively. The company has a tax rate of 35% and a cost of capital of 11.00%. What is the Economic value added (EVA)?

a.)497million

b.)625million

c.)460million

d.)350million
Answer #1

Answer : a)497 Million

EVA = NOPAT - (Capital Invested X Cost of Capital)

NOPAT = Earnings Before Interest and Taxes x (1-tax rate)

=1780 x(1-.35)

=1157

Capital Invested X Cost of Capital =( Long Term Debt + Share Holders equity) x 11%

=(2000+4000) x11% =660

EVA = NOPAT - (Capital Invested X Cost of Capital)

= 1157 -660

=497